# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v2

@@ 17 Receiver:
22 block_on_semaphore();//if semaphore==0, wait
@@
Figure 8: Flow of the PnM-based Covert Channel Attack.
@@
Sender-Receiver Synchronization. The sender and the receiver synchronize in order to ensure that the sender’s and receiver’s actions (i) are accurately aligned so that they do not interfere with each other and fail to communicate the message and (ii) to overlap the sender and receiver operations to increase the throughput of the attack. To achieve this, the sender and the receiver communicate using semaphores. The value of the semaphore indicates the number of batches that the sender has completed transmitting but the receiver has not yet completed probing. When the sender completes the transmission of a batch of bits, they increment the semaphore to signal the receiver to start probing the DRAM banks. This way, the sender overlaps the transmission of multiple batches of bits with the receiver’s probing operations without the need to wait for the receiver to complete the probing of the DRAM banks. Doing so, the sender can freely proceed to future batches without relying on synchronizing with the receiver. The receiver blocks on the semaphore until the sender increments it. When the semaphore is incremented, the receiver decrements it and starts probing the DRAM banks. When the receiver completes the probing of the DRAM banks, they read the semaphore again and the process repeats until the entire message is transmitted. In our attack implementation (see §5.1) we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
4.1.2 PuM-based Covert Channel
@@
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [55] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. [^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well. However, in this attack, we assume a PuM-based RowClone operation. For our attack, we focus on the Fast Parallel Mode in RowClone, which copies the data of a row to another row within a DRAM subarray.
@@
Figure 9 demonstrates how the content of the source row, i.e., src, is copied to the destination row, i.e., dst using RowClone’s Fast Parallel Mode. Initially, the rows and the row buffer are closed ( 1). RowClone operation consists of two key steps. First, the src is activated to store the entire row’s data in the row buffer ( 2). Second, the dst is activated, which leads to copying the contents of the row buffer in the dst ( 3). RowClone significantly reduces the latency and energy consumption of the bulk data copy operation from one row to another as it enables data movement without transferring data between the DRAM and the processing units. To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
Figure 9: Overview of RowClone’s Fast Parallel Mode.
@@
Attack Overview. PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied. Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks. The receiver repeatedly issues RowClone operations and measures the time it takes to complete these operations. This way, it can detect contention in a DRAM bank when another process accesses a different row in the same DRAM bank. If the sender issues a RowClone operation to a DRAM bank, it induces contention and the receiver detects a higher latency for its RowClone operation. This latency indicates that the corresponding bit is a logic-1. Otherwise, if the sender does not induce contention to a DRAM bank, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
@@
Figure 10 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
@@
Step 1. First, the receiver initializes all DRAM banks by issuing a RowClone operation to copy the contents of a source row to a destination row (Line 16). Then, the sender and the receiver synchronize on a barrier (Lines 10&17) which ensures that they start the communication turn at the same time.
@@
Step 2. The sender uses a single RowClone operation to transmit in parallel an N-bit message (where N is the number of banks) to the receiver and encodes the message in the mask operand of the RowClone operation (Lines 11-12). The sender encodes the message by setting the corresponding bits in the mask operand to 1 or 0 ( 1, 2). Doing so, the sender causes contention in the corresponding DRAM banks and activates a different row in the row buffer. To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4). The sender synchronizes with the receiver using a barrier operation (Lines 13&18).
Figure 10: End-to-End Flow of the Processing-using-Memory Covert Channel Attack.
@@ 15 Receiver:
Step 3. The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency. The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24). If the sender induces contention to a DRAM bank, the receiver high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26). Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
@@
Advantages over PnM-based Covert Channel. IMPACT-PuM improves upon the PnM-based covert-channel attack in a major way. Using a single RowClone operation, the sender can transmit in parallel an N-bit message to the receiver, where N is the number of DRAM banks which requires less computational resources to transmit the same amount of information compared to the PnM-based covert-channel attack. For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel while in the PnM-based covert-channel attack, the sender needs to separately execute a PEI instruction to activate each DRAM bank.
4.2 PnM-based Side-Channel Attack on DNA Read Mapping
@@
DNA Read Mapping. DNA is one of the most unique identifiers of an individual, such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity[108]. Therefore, preserving DNA privacy carries importance. DNA sequencing of a biological sample produces short DNA fragments called reads. Read mapping is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template genome to construct the complete genome of the sample.
@@
Figure 11 shows a high-level overview of a DNA read mapping (DNA-RM) workflow. DNA-RM has two major key steps: (i) seeding and (ii) alignment. During the seeding step, we identify the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome. After seeding, we align the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching location among all possible locations in the alignment step.
Figure 11: A simple overview of DNA read mapping workflow: (a) seeding and (b) alignment
@@
DNA-RM requires a large amount of data movement to access different parts of the hash table and then corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of DNA-RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 109, 110, 111, 112, 113, 114, 115]. In this work, we demonstrate that IMPACT can be used to launch a high-throughput side-channel attack against DNA-RM and quickly leak a large amount of information about the sample genome.
@@
Attack Overview. We assume that the victim is running a DNA-RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [116]), and the hash table is shared between the attacker and the victim. [^3] The hash table that contains information about the location of seeds in the reference genome is created once (acts as a database), and each user (attacker and victim in our case) of minimap2 probes it with queries. The attacker’s goal is to leak information about the test genome by observing the DNA-RM application’s memory access patterns.
@@
Figure 12 shows the high-level overview of the attack. The victim application hashes each seed of the test genome and offloads the corresponding seeding and alignment step to the PiM-enabled system  1. During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2. The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3. If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the test genome in the reference genome. As shown in previous works on DNA imputation [117, 118], the attacker can use the leaked information to infer properties about some regions of the private sample genome  4.
@@
Figure 12: High-level overview of IMPACT attack model on PiM-enabled DNA-RM
@@ 5.1 Methodology
Covert Channel Threat Model. Our threat model for the covert-channel attack envisions a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to user space programs without any restrictions and are not privileged instructions. The receiver is assumed to have access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access times. For IMPACT-PnM, the sender and the receiver need to colocate data in the same DRAM banks. For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity - when the sender issues a RowClone operations that span across multiple banks, all individual bank-level RowClone operations need to be completed before another process accesses DRAM.
@@
Side Channel Threat Model. Our threat model for the side-channel attack envisions a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use a DNA-RM implementation based on minimap2 [116] and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system. We assume that the alignment step includes chaining. We use the human genome as the reference and compare it against synthetic sample genomes. We assume that the hash table is distributed across each bank (i.e., DRAM employs a page interleaving mapping). We experiment with multiple different seed sizes, and we report the best-performing selection.
@@
Proof-of-Concept. We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [119]. We choose to construct the PoC using Sniper for two reasons. First, it is validated against modern Intel processors [120]. Second, there is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions in order to measure execution cycles from the user space.
@@
Table 2 shows the configuration our simulation setup. Before launching the attack, the sender and the receiver threads warm up by executing both attacks in order to avoid compulsory DRAM accesses for instructions or page table walks (i.e., after an L2 TLB miss, the page table walker might issue requests to DRAM) and reduce the noise of the attack. For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an addition operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
@@
Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of five different covert-channel attacks.
@@
(i) DRAMA-clflush: The main memory covert-channel attack proposed in [70] which uses clflush instructions to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
@@
(ii) DRAMA-Eviction: The main memory covert channel attack proposed in [70] which uses cache eviction sets to send memory requests directly to main memory. In both DRAMA attacks, the sender creates contention in the DRAM banks by causing row buffer conflicts and the receiver detects the contention to decode the message.
@@
(iii) Streamline: A state-of-the-art cache covert channel proposed in [121] which enables a flush-less collusion between the sender and the receiver.
@@
(iv) DMA Engine: A main memory attack that uses the DMA engine to bypass the caches and send memory requests directly to main memory. In this scenario, we assume a powerful attacker can infer the DRAM access time directly from the user space by disregarding context switch and OS-related latencies.
@@
(v) PnM-OffChip: A PnM-based attack that probes the locality monitor and a state-of-the-art off-chip predictor [122] before offloading PEI instructions to compute units near DRAM banks. If the off-chip predictor predicts that data experience high locality, the computation is performed by the CPU instead of the PnM, leading to reduced throughput.
@@
(iv) IMPACT-PnM: Our proposed attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel between the sender and the receiver.
@@
(v) IMPACT-PuM: Our proposed attack that makes use of RowClone to offload bulk-copy operations directly to main memory and establish a covert channel between the sender and the receiver.
@@
For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack. To evaluate the throughput of the DRAMA and Streamline-based attacks, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack. We validated the analytical model against the throughput achieved by DRAMA and Streamline in a real system. For example, in Streamline, the authors report a maximum communication throughput of 1.8M Mb/s and our model estimates that the upper bound of the throughput is 2.7 Mb/s for the smallest LLC size.
Noise Sources in Simulation Setup and Error Rate. There are two main microarchitectural structures that generate noise in the simulated system: 1) hardware prefetchers and 2) page table walkers. To observe the impact of application-induced noise on the throughput of IMPACT attacks, we run applications from GraphBIG [123] benchmark suites concurrently with attacker and victim (or sender and receiver) processes. We measure the throughput of each attack by incorporating the error rate in the measurement: we measure the throughput only based on successful leakage.
@@
MMU L1 I-TLB: 128-entry, L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access, 3 Split PWCs: 32-entry, 2-cycle access

### v3

@@ 17 Receiver:
22 block_on_semaphore();
23 //if semaphore==0, wait
@@
Step 1. The receiver issues PEIs to initialize each DRAM bank by activating a predetermined row (Line 18). The receiver either uses unused cache lines of the activated row or sets the ignore flag in the corresponding entry in order to bypass the locality monitor. Then, the sender and the receiver synchronize on a barrier (Lines 5 and 20), which ensures that they start the communication turn at the same time.
Step 2. The sender starts the communication turn by issuing PEI instructions to activate a row in each DRAM bank (Lines 7-12) or not based on the bit value that will be transmitted. To transmit a 0 ( 1), the sender does not activate any row in the corresponding DRAM bank to avoid causing interference in the row buffer (e.g., Row 0xB remains in the row buffer  2). To transmit a 1 ( 3), the sender activates a different row to cause interference in the corresponding DRAM bank (i.e., a row buffer conflict between Row 0xD and Row 0xE  4). At the end of the communication turn (after transmitting M bits), the sender first executes a memory fence (Line 14) to ensure that all the PEI instructions have been executed and then signals the receiver to start probing the DRAM banks (Line 15).
Step 3. The receiver starts probing the DRAM banks by issuing PEI instructions to activate the same row as the one activated during initialization (Line 25). Then, the receiver measures the time it takes to complete the request (Line 27). If the latency exceeds a predetermined hit threshold, the receiver senses the interference in the row buffer and decodes the bit as a 1 (Lines 28-29  5, 6). Otherwise, the access is considered a row buffer hit and the receiver decodes the bit as a 0 (Lines 30-31  7, 8). The sender and the receiver repeat the same steps for each batch until the entire message is transmitted.
@@
Sender-Receiver Synchronization. The sender and the receiver synchronize to 1) ensure their actions do not interfere with each other and 2) overlap the latencies of their operations to increase the throughput of the attack. To do so, the sender and the receiver use a semaphore. The value of the semaphore indicates the number of batches that the sender has completed transmitting, but the receiver has not yet probed. The receiver blocks on the semaphore until the sender increments it (i.e., does not start probing). When the sender completes the transmission of a batch, they increment the semaphore to signal the receiver to start. Upon receiving the signal, the receiver decrements the semaphore and starts probing the DRAM banks. In parallel, the sender can start transmitting the next batch that utilizes a different set of DRAM banks. This way, they can overlap the transmission latency and the probing latency. When the receiver completes probing, they read the semaphore again, and the process repeats until the entire message is transmitted. In our attack implementation (§5.1), we use POSIX semaphores to implement the sender-receiver synchronization and use a batch size of 4 bits.
4.2 PuM-based IMPACT Covert-Channel Attack
@@
Baseline PuM Architecture. We assume a PuM architecture that provides user applications with RowClone [52] that enables bulk data movement operations (e.g., bulk data copy and initialization) within DRAM’s subarray boundaries by exploiting the internal operations of DRAM. [^2] There could be near-memory RowClone variants that move data without using operational principles of DRAM as well. However, in this attack, we assume a PuM-based RowClone operation. RowClone’s Fast Parallel Mode works by issuing back-to-back two activation commands: the first activation copies the content of a source row (i.e., src) into the row buffer. The second command connects the destination row (i.e., dst) to the bitlines. Since the row buffer already stored the src’s data, by the time the dst is activated, the data in dst is overwritten by src’s data.
@@
To execute RowClone, we assume that the user space application specifies a source and a destination virtual address range and the corresponding mask to copy only specific parts of the source to the destination. This way, the memory controller receives one RowClone request and breaks it into parallel RowClone requests, one for each set bit of the mask.
@@
Attack Overview. PuM-based covert channel attack (IMPACT-PuM) exploits in-DRAM bulk copy operations to establish a covert channel between a sender and a receiver. The sender transmits a message by executing a RowClone operation with three operands: (i) a source address range, (ii) a destination address range, and (iii) a mask to select which parts of the ranges will be copied. Source and address ranges correspond to physical pages spanning across all DRAM banks and by using the mask, the sender can encode the transmitted message by causing contention only in selected DRAM banks. The receiver repeatedly issues RowClone operations and measures the latency of these operations. This way, if the sender issues a RowClone operation to a DRAM bank, it induces contention, and the receiver detects a higher latency for its RowClone operation, indicating the transmitted bit is logic-1. Otherwise, the receiver detects a lower latency for the RowClone operation, indicating that the corresponding bit is a logic-0.
@@
Fig. 5 and Listing 2 show the end-to-end flow of IMPACT-PuM and all the steps involved to establish the covert channel between the sender and the receiver.
Figure 5: PuM covert channel attack flow.
@@
Step 1. First, the receiver initializes all DRAM banks by issuing a RowClone operation (Line 16). Then, the sender and the receiver synchronize on a barrier (Lines 10 and 17) which ensures that they start the communication turn at the same time.
@@
Step 2. The sender uses a single RowClone operation to transmit an N-bit message (where N is the number of banks) to the receiver in parallel and uses the mask of the RowClone operation to encode the message (Lines 11-12). The sender encodes the message by setting the corresponding bits in the mask to 1 or 0 ( 1, 2). To send a 1, the sender triggers RowClone in the corresponding DRAM bank and induces contention by altering the activated row in the row buffer ( 3). To send a 0, the sender does not trigger RowClone in the corresponding DRAM bank and does not induce contention ( 4). The sender synchronizes with the receiver using a barrier operation (Lines 13 and 18).
@@ 15 Receiver:
Step 3. The receiver in IMPACT-PuM detects contention in DRAM banks by issuing RowClone operations and measuring their latency. The receiver continuously issues one RowClone operation for each DRAM bank (in contrast to the sender who issues only a single RowClone operation), to the initial virtual address range (i.e., rows initialized in Step 1) and measures the latency of each operation ( 2, Line 24). If the sender induces contention in a DRAM bank, the receiver observes high latency for the RowClone operation which indicates that the corresponding bit in the message is set to 1 ( 3, Lines 25-26). Otherwise, the receiver detects a fast RowClone operation which indicates that the bit is set to 0 ( 4, Lines 27-28).
@@
Advantage over IMPACT-PnM. In IMPACT-PuM, the sender can use a single RowClone operation to transmit an N-bit message to the receiver in parallel, where N is the number of DRAM banks. Thus, IMPACT-PuM requires less computational resources to transmit the same amount of information compared to IMPACT-PnM. For example, using a single thread, the sender can activate all N DRAM banks and transmit N bits of information in parallel. In contrast, in IMPACT-PnM, the sender needs to execute one PEI per bit.
4.3 PnM-based Side-Channel Attack on Read Mapping
We build a PnM-based side-channel attack that exploits PiM operations to leak private information of a concurrently running application. To demonstrate the IMPACT side-channel attack, we target a DNA sequence analysis application.
@@ 4.3 PnM-based Side-Channel Attack on Read Mapping
Read Mapping. DNA is one of the unique identifiers of an individual [73], such that even a small fraction of DNA can include sufficient information for linking to the owner’s identity [70, 71]. Therefore, preserving privacy while processing genomic data carries utmost importance to protect sensitive information. Sequencing the DNA of a biological sample produces short DNA fragments called reads. The reads are then processed by read mapping, which is a fundamental operation to find the matching locations of reads in the reference genome that acts as a template to construct the complete genome of the sample.
@@
Read mapping (RM) has two key steps: (i) seeding and (ii) alignment. During the seeding step, RM identifies the possible locations of the reads on the reference genome based on sequence similarities between the reads and the reference genome. Seeding involves hashing small segments of the DNA sequences (i.e., seeds) and probing a hash table that is built from the reference genome. After seeding, RM aligns the reads with their candidate regions on the reference (i.e., potential reference segments) using dynamic programming-based methods to find each read’s best matching position among all possible locations in the alignment step.
@@
RM requires a large amount of data movement to access different parts of the hash table and corresponding regions on the reference genome. Hence, several prior works aim to improve the performance of RM by offloading seeding and alignment to PiM-enabled systems to reduce the data movement overheads [33, 96, 97, 98, 99, 100, 101, 102]. In this work, we demonstrate that an attacker can launch a high-throughput side-channel attack against RM by utilizing PiM instructions and quickly leak a large amount of information about the sample genome.
@@
Attack Overview. We assume that the victim is running a RM application on a PiM-enabled system and the attacker is able to run a malicious application on the same system. The attacker and the victim are using the same read mapping tool (e.g., minimap2 [103]). The read mapping tool constructs a hash table that contains information about the seed locations in the reference genome, and allows each user in the system to probe the hash table via queries during read mapping. We assume the hash table is distributed across multiple DRAM banks. This assumption is realistic as many modern DRAM address mapping schemes (e.g., [104, 105, 106, 107]) aim to interleave consequent memory chunks across different DRAM banks to exploit bank-level parallelism [108, 109].
@@
Fig. 6 shows the high-level overview of the attack. The victim application extracts the seeds from the reads and offloads the corresponding seeding and alignment step to the PiM-enabled system  1. During the seeding step, the corresponding DRAM row gets activated (i.e., Index 0-5 of the hash table)  2. The attacker continuously uses PnM instructions to probe the shared hash table in an attempt to identify if the victim application is accessing a specific index of the hash table  3. If the attacker observes that the victim application is accessing a specific index, it retrieves partial or exact information about the potential locations of the reads in the reference genome. As shown in previous works on DNA imputation [110, 111, 112, 113], the attacker can use the leaked information in a completion attack [71] to infer properties about some regions of the private sample genome  4.
@@
Figure 6: High-level overview of IMPACT attack model on PiM-enabled RM.
@@ 5.1 Methodology
Covert Channel Threat Model. We assume a scenario where a sender and a receiver operate on the same PiM-enabled system to exchange information. PiM instructions are available to user space applications without any restrictions. The receiver has access to cpuid and rdtscp instructions, enabling high-precision measurement of memory access latencies. For IMPACT-PnM, the sender and the receiver co-locate data in the same DRAM banks. For IMPACT-PuM, (i) the sender allocates two virtual address ranges whose corresponding physical pages span across all DRAM banks, and (ii) RowClone guarantees atomicity: when the sender issues a RowClone operation that spans across multiple banks, all individual bank-level RowClone operations need to be completed before another DRAM operation is executed.
@@
Side Channel Threat Model. We assume a scenario where an attacker process operates on the same PiM-enabled system as a victim process and leaks information about the victim process. We use an RM implementation based on minimap2 [103], and we modify the implementation to offload the seeding and alignment step to the PiM-enabled system. We assume that the alignment step includes chaining. We use the human genome as the reference and compare it against synthetic sample genomes. We experiment with multiple seed sizes and report the best-performing selection.
@@
Proof-of-Concept. We construct a proof-of-concept of all IMPACT variants using the Sniper Multicore Simulator [114]. [^3] There is no commercially available system that follows a similar architecture as PEI or any modern system that enables end-to-end RowClone operations. We extend Sniper to (i) accurately model all internal DRAM operations, (ii) emulate and simulate the PiM interface and PiM operations, and (iii) emulate the functionality of the cpuid and the rdtscp instructions to measure execution cycles from the user space.
@@
Table 2 shows the configuration of our simulation setup. Before launching the attack, the sender and the receiver threads warm up by executing both attacks to avoid compulsory DRAM accesses for instructions or page table walks. For IMPACT-PnM, we assume that both the attacker and the receiver use PEIs to perform an operation that takes $\sim 3$ cycles to execute (in addition to the actual DRAM access).
@@
Comparison Points for Covert-Channel Attacks. To showcase the effectiveness of IMPACT, we evaluate the leakage throughput of seven different covert-channel attacks.
@@
(i) DRAMA-clflush: The row buffer-based covert-channel attack proposed in [68] which uses clflush instructions to send memory requests directly to main memory. clflush only probes the LLC to flush the cache line.
@@
(ii) DRAMA-Eviction: The row buffer-based covert channel attack proposed in [68] which uses cache eviction sets to send memory requests directly to main memory.
@@
(iii) Streamline: State-of-the-art cache covert channel proposed in [115] which enables a flushless communication.
@@
(iv) DMA Engine: A row buffer-based attack that uses the DMA engine to bypass the caches. We assume a powerful attacker can infer the DRAM access time directly from the user space by ignoring context switch and OS-related latencies.
@@
(v) PnM-OffChip: A PnM-based attack in a PnM architecture with locality monitor and a state-of-the-art off-chip predictor [116]. If the data is predicted to have high locality, the computation is performed by the CPU instead of the near-memory compute units, leading to reduced attack throughput.
@@
(vi) IMPACT-PnM (§4.1): An attack that bypasses the locality monitor and offloads PEI instructions directly to compute units near DRAM banks in order to establish a covert channel.
@@
(vii) IMPACT-PuM (§4.2): An attack that leverages RowClone to offload bulk-copy operations directly to main memory and establish a covert channel.
@@
For a fair comparison against DRAMA and Streamline, we showcase the upper bound of the communication throughput achieved by each attack. To calculate their throughput, we use our simulation infrastructure to extract parameters such as the LLC hit latency, average LLC miss latency, cache lookup latency, cache hit/miss ratio, and feed them in an analytical model that calculates the maximum throughput of the attack. We validate the model against the real system throughput of DRAMA and Streamline (e.g., Streamline [115] reports a maximum communication throughput of 1.8 Mb/s and we model the upper bound as 2.7 Mb/s for the smallest LLC).
Noise Sources in Simulation Setup and Error Rate. We simulate hardware prefetchers and page table walkers to induce noise in the simulated system. We measure the throughput of each attack only based on the successfully leaked data (i.e., incorporating the error rate in our measurements).
@@
MMU L1 DTLB (4KB): 64-entry, 4-way assoc, 1-cycle access, L1 DTLB (2MB): 32-entry, 4-way assoc, 1-cycle access, L2 TLB: 1536-entry, 12-way assoc, 12-cycle access